Calculate the maximum monochromatic emissive power at 288k
1
Expert's answer
2019-05-27T11:29:29-0400

Wien’s Displacement Law:

"E_{\lambda}=\frac{C_1\lambda^{-5}}{exp^{\frac{C_2}{\lambda T}}-1}"

the wavelength at which emissive power is maximum:


"\lambda_{max}=\frac{C_2}{4.965\times T}=\frac{0.0029}{T} m"

If we substitute this wavelength in the Wien's Displacement law, then the magnitude of the maximum monochromatic emissive power is


"E_{max}=1.285\times 10^{-5}\times T^5=\\
1.285\times 10^{-5}\times 288^5 = 25.46\times 10^6 \frac{W}{m^2}"